A new endocrine cancer diagnostic algorithm is tested on 70 subjects. 10 of them are actual cancer patients and 60 of them are cancer-free. The diagnostic algorithm has 90% sensitivity and 80% specificity on these subjects. What are the numbers of True Negative and False Positive cases?
Answer #1

numbers of True Negative =P(tested negative given not have disease)

=N(not have disease)* specificity =(70-10)*0.8 =48

numbers of False Positive =N(not have disease)-numbers of True Negative=60-48 =12